IfA+B=135find. ,(1+cotA)(1+cotB).??? give me the answer please

Grade:11

1 Answers

poojitha
6 years ago
A+B=135,cot(A+B)=cot135,(cotB-cotA)/(1+cotA.cotB)=-1, 1+cotA.cotB=cotA-cotB ,                                        (1+cotA).(1+cotB)=1+cotA+cotB+cotA.cotB=(1+cotA.cotB)+cotA+cotB=cotA-cotB+cotA+cotB=2cotA